A charge ݱQ is placed on the plates of a capacitor, which is then disconnected from the outside world. The separation between the plates is changed, and it is found that the voltage across the plates decreases.

(a) Have the plates been moved closer together or farther apart?
closer together
farther apart
(b) If the new voltage is smaller than the original voltage by a factor of four, what is the ratio of the new plate spacing to the original spacing?
Answer #1

a.) closed together as V = Qd/Ae

b.) d1/d1 = V1/V2

=> v2 = V1/4